# BihoT: A Large-Scale Dataset and Benchmark for Hyperspectral Camouflaged Object Tracking

### v1

@@ Abstract
Hyperspectral object tracking (HOT) has exhibited potential in various applications, particularly in scenes where objects are camouflaged. Existing trackers can effectively retrieve objects via band regrouping because of the bias in existing HOT datasets, where most objects tend to have distinguishing visual appearances rather than spectral characteristics. This bias allows the tracker to directly use the visual features obtained from the false-color images generated by hyperspectral images without the need to extract spectral features. To tackle this bias, we find that the tracker should focus on the spectral information when object appearance is unreliable. Thus, we provide a new task called hyperspectral camouflaged object tracking (HCOT) and meticulously construct a large-scale HCOT dataset, termed BihoT, which consists of 41,912 hyperspectral images covering 49 video sequences. The dataset covers various artificial camouflage scenes where objects have similar appearances, diverse spectrums, and frequent occlusion, making it a very challenging dataset for HCOT. Besides, a simple but effective baseline model, named spectral prompt-based distractor-aware network (SPDAN), is proposed, comprising a spectral embedding network (SEN), a spectral prompt-based backbone network (SPBN), and a distractor-aware module (DAM). Specifically, the SEN extracts spectral-spatial features via 3-D and 2-D convolutions to form a refined prompt representation. Then, the SPBN fine-tunes powerful RGB trackers with spectral prompts and alleviates the insufficiency of training samples. Moreover, the DAM utilizes a novel statistic to capture the distractor caused by occlusion from objects and background and corrects the deterioration of the tracking performance via a novel motion predictor. Extensive experiments demonstrate that our proposed SPDAN achieves state-of-the-art performance on the proposed BihoT and other HOT datasets.
@@
Visual object tracking [1, 2] is a crucial task that involves finding the same objects in subsequent frames based on their initial positions in the first frame. It has great potential in various applications such as autonomous driving, video analysis, and so on [3, 4, 5]. However, visual object tracking makes it difficult to handle interference caused by similar appearances and background clutter. Hyperspectral images (HSIs) contain the texture and spectral information of an object and can use spectral information to track an object in the presence of visual interference. Hence, hyperspectral object tracking (HOT) has attracted more attention from the research community.
@@
With the construction of HOT datasets such as HOT-2020 [6], IMEC-25 [7], and WHU-Hi-H3 [8], many deep learning-based methods have been proposed [9, 10, 11, 12, 13]. Although the above methods have achieved great progress, many discriminative spectral features still have not been explored. As shown in Fig. 1(a), we argue that the existing HOT dataset has two characteristics: one is the visual distinguishable (v-dis) factor, which means that using visual information can correctly track objects, and the other is the spectral distinguishable (s-dis) factor, which means that using spectral information can correctly track objects. Taking the HOTC-2020 dataset as an example, video sequences with the v-dis factor account for 77.1% of the entire dataset, while the s-dis factor only accounts for 22.9% of the entire dataset. Considering that most existing HS tracking methods rely on pre-trained RGB trackers for processing false-color images generated from HS images [10, 11, 14], which are good at handling v-dis factor. So in many cases, even if the tracker cannot extract robust spectral features, it can still capture objects through visual information (v-dis factor). This bias of datasets makes most algorithms overlook the utilization of the advantages of hyperspectral information.
@@
Despite existing HOT tasks exploring the degradation of object visual information due to factors like illumination changes and motion blur, objects are typically in a v-dis state. This condition does not make the trackers leverage the spectral characteristics of objects, thereby limiting their ability to handle tracking tasks in complex environments. To address this problem, we introduce a novel hyperspectral camouflaged object tracking (HCOT) task, which focuses on tracking objects that share the same visual information but differ in spectral information. Considering that no existing dataset suits this task, we construct the first large-scale HCOT dataset, BihoT. This dataset has two significant advantages over existing datasets. First, BihoT collects many visually similar objects, called camouflaged objects. As illustrated in Fig. 1(b), the plastic fake fruit is a camouflaged object relative to the real fruit. The proportion of the s-dis factor reaches 95.5%, while the proportion of the v-dis factor is only 4.5%. For camouflaged objects, the v-dis factor is usually unreliable in tracking, underscoring the importance of the spectral (s-dis) factor. Second, BihoT consists of 49 hyperspectral video (HSV) sequences, containing 41,912 HSI in total, making it the largest HCOT dataset with 25 bands. Each HS image encompasses 25 bands and a resolution of 409 217, meeting the data requirements for deep learning.
@@
Furthermore, to promote the development of HCOT, we propose a spectral prompt-based distractor-aware network (SPDAN) as a simple yet effective baseline to tackle the challenges of spectral feature extraction and frequent occlusion. Specifically, a spectral embedding network (SEN) is designed to transform the spatial and spectral information to the tokenized features, which offers discriminative spectral information for existing patch embedding methods [15] via a 3-D convolutional layer. Then, motivated by prompt learning [16], a spectral prompt-based backbone network (SPBN) is proposed to transfer the knowledge from the RGB domain to the HS domain to extract the robust spectral-aware features. Finally, to address the challenges of frequent occlusion in practical applications, we propose a distractor-aware module (DAM) to capture frames with disturbances such as occlusion accurately. Different from the SENSE [10], DAM considers both APCE [17] and the offset of the prediction boxes across frames, which can make more accurate judgments on distractors and adopt motion state models to predict states with occlusion, effectively mitigating the decline in the tracking performance caused by such distractors. As shown in Fig. 2, it can be observed that our proposed method outperforms the other state-of-the-art HOT methods on the HOTC-2020 dataset. The main contributions of this paper are as follows:
@@
The first large-scale HCOT dataset, BihoT, is constructed, which contains 41,912 HS camouflaged images paired with the bounding box annotations. It provides more challenging scenes and brings profound insights for HOT.
@@
A novel spectral embedding network is designed to transfer the spatial and spectral information to the tokenized features, and aggregate the spectral and visual features from the prompt-based learning to strengthen camouflaged feature representation.
@@
A distractor-aware module is proposed to perceive the distractor by considering both the average peak-to-correlation energy and the offset of the predicted boxes, and rectify the tracking results through a motion predictor, improving the robustness of the tracker to distractors.
The remaining structure of the paper is structured as follows. Section II reviews some related works. Section III introduces the proposed BihoT dataset. Section IV formulated the proposed methods. Section V provides the ablation and comparative experimental results. Finally, conclusions are drawn in Section VI.
@@ II-A HS Object Tracker
HOT aims to discover the same objects in subsequent frames based on the initial object state and spectral information. Early work mostly used traditional methods for predicting the object position, such as histogram feature-based methods [6, 19], correlation filters-based methods [20, 21, 22], and spectral reflectance-based methods [23], which has a limited ability to deal with the complicated distractors.
@@
Many deep learning-based methods have been recently proposed to extract robust deep features and handle complex scenes such as occlusion. It can be mainly categorized into band selection-based methods and feature fusion-based methods. The former methods leverage the RGB tracker to predict the motion of objects via the false-color images generated from HSIs. For example, Li et al. [14] proposed an ensemble learning-based method based on the false-color images generated by a band-selection method. The network produced the band weights via an auto-encoder-like structure, which is used to regroup multiple bands to several three-channel false-color images. Li et al. [24] designed an attention-based band grouping method to generate the importance weights of different bands. Chen et al. [10] proposed a spectral-spatial self-expression module to transform the HSI into false-color images by filling the modality gap. Chen et al. [11] presented a spectral awareness module to explore the band weight distributions by the multi-level band interactions. Gao et al. [25] manually selected the different bands to form the false-color images and extracted the bidirectional fused features to construct the robust features. The latter methods mainly adopt the Siamese tracker framework, comprising the spectral and RGB tracking branches to mine the cross-modality semantic cues. For example, Liu et al. [12] presented a spectral and visual feature fusion framework SiamHYPER to provide a rich semantic feature representation and obtained the accurate tracking results from the cross-modality complementary decision. Zhao et al. [26] fused features from spectral and RGB domains by the Transformer module to capture the correlation between the template images (TI) and search images (SI). Wu et al. [9] proposed a Siamese Transformer tracker to learn the heterogeneous features from different HSIs by the domain label reverse learning. Jiang et al. [27] proposed a unified network, termed as SiamCAT, to process HSIs that have varied numbers of bands. Besides, with a guided learning attention module, the spectral features can be highlighted by the spatial attention obtained from the RGB branch. Li et al. [28] designed a material-guided multiview fusion network MMF-Net, which combines the RGB, spectral, and material information to form a unified rich representation of an object.
@@ II-B HOT Datasets
It can be seen that the above datasets neglect the camouflage characteristics of objects and only collect images in natural scenes, with most objects being vehicles and pedestrians, lacking the challenge of visual similarity. Although the WHU-Hi-H3 dataset considers the similarity between objects relative to the background, it does not highlight the visual confusion caused by the similarity between objects. Therefore, to facilitate research on HCOT, we construct a large-scale 25-band dataset, BihoT, which contains 41,912 frames. The main shooting objects include fruits and daily necessities.
II-C Transformer-based Object Tracker
@@
In recent years, the Transformer has been applied in many computer vision tasks such as detection, segmentation, and classification due to its outstanding ability to model long-term dependencies. Many Transformer-based tracking methods have been proposed. Xie et al. [29] proposed a novel object-dependent feature network to embed the inter-image correlation. The non-object features can be suppressed in the self- and cross-attention schemes. Song et al. [30] designed a multi-scale cyclic shifting window attention to capture the window-level features. Mayer [31] proposed a Transformer-based model prediction module, which captures global correlations without inductive bias. Blatter et al. [32] designed an exemplar Transformer using the instance-level exemplar attention with the refined scaled dot-product calculation. Gao et al. [33] proposed an attention in attention module, which can find the consensus among all correlation vectors to enhance correlations. Cui et al. [34] presented a compact tracking framework MixFormer, which extracts and fuses features simultaneously through a mixed attention module without the need for additional fusion modules or box post-processing. Zhu et al. [16] developed a visual prompt tracking framework to achieve task-oriented multi-modality tracking. With the guidance of learned prompts, existing RGB-domain foundation models can effectively handle the downstream multi-modality tracking tasks. For HS trackers, there is currently limited research based on transformers. Gao et al. proposed a novel transformer-based tracking algorithm (Trans-HST) with a cross-band feature enhancement module to reduce the interaction information loss.
@@ III-A Image Collection
We employ an IMEC XIMEA MQ022HG-IM-SM5X5-NIR camera and an EDMUND OPTICS 35mm C VIS-NIR lens to capture the high spatial and spectral resolution HS images across the wavelength range of 600-974 nm. The original image is a two-dimensional image with pixels of . After processing with the transformation algorithm, a 3D hyperspectral cube of can be obtained. To ensure the diversity of the scene, we collect data from different seasons and illuminations. The objects are real and artificial plastic fruits or vegetables in daily life to ensure the acquisition of camouflaged objects. which is shown in Fig. 3
III-B Data Annotation and Condition Challenges
The BihoT dataset has 49 annotated sequences, of which 27 are for training and 22 are for testing. Each image has an object marked by a bounding box. All bounding boxes were manually created using the Labelme toolkit. In addition, we set 9 attributes corresponding to different challenging factors in the testing set, including Background Clutter (BC), Fast Motion (FM), In-Plane Rotation (IPR), Illumination Variation (IV), Low Resolution (LR), Occlusion (OCC), Out-of-Plane Rotation (OPR), Spectral Consistency (SC), and Spectral Variation (SV). Table II lists the number of bounding boxes and challenging scenarios for each sequence. To measure the spectral feature extraction ability of the tracker, we proposed two challenging factors, Spectral Consistency (SC) and Spectral Variation (SV), compared with the existing datasets [6, 8, 7]. Specifically, in the BihoT dataset settings, there are multiple real objects and one camouflaged object in each sequence. SC means that when the real object moves, its spectral information remains unchanged relative to other real objects in the scene. When the camouflaged object moves, the spectral information relative to other real objects in the scene is different. Therefore, the scene where the camouflaged object moves is denoted as SV.

### v2

@@ Abstract
Hyperspectral object tracking (HOT) has many important applications, particularly in scenes where objects are camouflaged. The existing trackers can effectively retrieve objects via band regrouping because of the bias in the existing HOT datasets, where most objects tend to have distinguishing visual appearances rather than spectral characteristics. This bias allows a tracker to directly use the visual features obtained from the false-color images generated by hyperspectral images without extracting spectral features. To tackle this bias, the tracker should focus on the spectral information when object appearance is unreliable. Thus, we provide a new task called hyperspectral camouflaged object tracking (HCOT) and meticulously construct a large-scale HCOT dataset, BihoT, consisting of 41,912 hyperspectral images covering 49 video sequences. The dataset covers various artificial camouflage scenes where objects have similar appearances, diverse spectrums, and frequent occlusion, making it a challenging dataset for HCOT. Besides, a simple but effective baseline model, named spectral prompt-based distractor-aware network (SPDAN), is proposed, comprising a spectral embedding network (SEN), a spectral prompt-based backbone network (SPBN), and a distractor-aware module (DAM). Specifically, the SEN extracts spectral-spatial features via 3-D and 2-D convolutions to form a refined prompt representation. Then, the SPBN fine-tunes powerful RGB trackers with spectral prompts and alleviates the insufficiency of training samples. Moreover, the DAM utilizes a novel statistic to capture the distractor caused by occlusion from objects and background, and corrects the deterioration of the tracking performance via a novel motion predictor. Extensive experiments demonstrate that our proposed SPDAN achieves state-of-the-art performance on the proposed BihoT and other HOT datasets.
@@
Visual object tracking [1, 2] is a crucial task in many applications, such as autonomous driving, video analysis, etc. [3, 4, 5]. It is to find the same objects across consecutive frames based on their initial positions in the first frame. However, it faces challenges with interference from similar appearances and background clutter. Hyperspectral images (HSIs), which provide both texture and spectral information, can effectively track objects despite visual interference, which is widely used in remote sensing fields [6, 7]. Thus, hyperspectral object tracking (HOT) has garnered increased attention from researchers.
@@
With the construction of HOT datasets such as HOT-2020 [8], IMEC-25 [9], and WHU-Hi-H3 [10], many deep learning-based methods have been proposed [11, 12, 13, 14, 15]. While these methods have demonstrated substantial progress, many discriminative spectral features remain under-explored. As illustrated in Fig. 1(a), we identify two key characteristics in the existing HOT datasets: 1) the visually distinguishable (v-dis) factor, which indicates that objects can be accurately tracked using visual information, and 2) the spectral distinguishable (s-dis) factor, which reflects the ability to track objects based on spectral information. Taking the HOTC-2020 dataset as an example, video sequences with the v-dis factor account for 77.1% of the entire dataset, while those with the s-dis factor only account for 22.9%. Considering that most existing HOT methods rely on pre-trained RGB trackers for processing false-color images generated from HS images [12, 13, 16], which are good at handling v-dis factor, even when a tracker struggles to extract reliable spectral features, it can still identify objects by leveraging visual cues (v-dis factor). This bias of datasets makes most existing algorithms overlook the advantages of hyperspectral (HS) information.
@@
Despite the existing HOT tasks exploring the degradation of object visual information due to factors like illumination changes and motion blur, objects typically remain in a v-dis state. This condition prevents trackers from fully leveraging the spectral characteristics of objects, thereby limiting their ability to handle complex tracking environments. To address this challenge, we introduce a novel task, Hyperspectral Camouflaged Object Tracking (HCOT), which focuses on tracking objects that share the same visual information but differ in spectral characteristics. This technology enhances the focus on tracked objects while mitigating interference from artificial decoys, thereby improving the accuracy of criminal tracking and the reliability of security systems. Considering that no existing dataset suits this task, we construct the first large-scale HCOT dataset, BihoT. This dataset offers two significant advantages over existing datasets. First, BihoT collects numerous visually similar objects—camouflaged objects. As illustrated in Fig. 1(b), the plastic fake fruit is a camouflaged object relative to the real fruit. In BihoT, the proportion of the s-dis factor reaches 95.5%, while the proportion of the v-dis factor is only 4.5%. For camouflaged objects, the v-dis factor is typically unreliable in tracking, underscoring the importance of the s-dis factor. Second, BihoT comprises 49 hyperspectral video sequences (HSVs) containing 41,912 HSIs in total, making it the largest HCOT dataset with 25 bands. Each HS image encompasses 25 bands and a resolution of 409 217, meeting the data requirements for deep learning.
@@
Furthermore, to promote the development of HCOT, we propose a spectral prompt-based distractor-aware network (SPDAN), a simple yet effective baseline designed to address the challenges of spectral feature extraction and frequent occlusion. Existing SOTA HS trackers merely adjust the number of input channels in existing models, such as ResNet, without effectively extracting fine-grained information from the spectral dimension [14, 11]. Thus, a spectral embedding network (SEN) is designed to transform the spatial and spectral information into tokenized features via a 3-D convolutional layer. The 3-D convolutions simultaneously establish spectral and spatial correlations across multiple bands, allowing it to capture pixel variations between different bands and extract rich spectral information [17, 18, 19]. Then, motivated by prompt learning [20], a spectral prompt-based backbone network (SPBN) is proposed to transfer knowledge from the RGB domain to the HS domain, enabling robust spectral feature extraction. To tackle frequent occlusions in practical applications, we introduce a distractor-aware module (DAM) to capture frames with disturbances such as occlusion accurately. Different from [21], DAM considers both APCE and the offset of the prediction boxes across frames, which can make more accurate judgments on distractors. Moreover, it incorporates motion state models to predict object states during occlusion, effectively mitigating the decline in tracking performance caused by such distractors. The main contributions of this paper are as follows:
@@
The first large-scale HCOT dataset, BihoT, is constructed, which contains 41,912 HS camouflaged images paired with the bounding box annotations. This dataset introduces more challenging scenarios and provides valuable insights for advancing HOT research.
@@
A novel spectral embedding network is designed to transfer the spatial and spectral information to the tokenized features. SEN effectively aggregates spectral and visual features through prompt-based learning, enhancing the representation of camouflaged objects.
@@
A distractor-aware module is proposed to perceive the distractor by considering both the average peak-to-correlation energy and the offset of the predicted bounding boxes. DAM rectifies tracking results using a motion predictor, improving the robustness against distractors.
The remaining of this paper is organized as follows. Section II reviews some related works. Section III introduces the proposed BihoT dataset. Section IV formulates the proposed methods. Section V provides the ablation and comparative experimental results. Finally, conclusions are drawn in Section VI.
@@ II-A HS Object Tracker
HOT aims to discover the same objects in subsequent frames based on the initial object state and spectral information. Early works mostly used the traditional methods for predicting the object position, such as histogram feature-based methods [8], correlation filters-based methods [23, 24], and spectral reflectance-based methods [25], which has a limited ability to deal with the complicated distractors.
@@
Many deep learning-based methods have been recently proposed to extract robust deep features and handle complex scenes such as occlusion. It can be mainly categorized into band selection-based methods and feature fusion-based methods. The former methods leverage the RGB tracker to predict the motion of objects via the false-color images generated from HSIs. For example, Li et al. [16] proposed an ensemble learning-based method based on the false-color images generated by a band-selection method. The network produced the band weights via an auto-encoder-like structure, which is used to regroup multiple bands to several three-channel false-color images. Li et al. [26] designed an attention-based band grouping method to generate the importance weights of different bands. Chen et al. [12] proposed a spectral-spatial self-expression module to transform the HSI into false-color images by filling the modality gap. Chen et al. [13] presented a spectral awareness module to explore the band weight distributions by the multi-level band interactions. Methods based on band selection can leverage the processing capabilities of RGB trackers to predict object positions from three-channel images. However, this is due to the inherent bias in existing datasets. The v-dis factor of HS objects collected in natural scenes tends to dominate, while the s-dis factor is often overlooked. As a result, the tracking algorithm is not effectively guided to explore the richer spectral information available in HS data, which fails to meet practical application requirements.
The latter methods mainly adopt the Siamese tracker framework, comprising the spectral and RGB tracking branches to mine the cross-modality semantic cues. For example, Liu et al. [14] presented a spectral and visual feature fusion framework SiamHYPER to provide a rich semantic feature representation and obtained accurate tracking results from the cross-modality complementary decision. Wu et al. [11] proposed a Siamese Transformer tracker to learn the heterogeneous features from different HSIs by the domain label reverse learning. Jiang et al. [27] proposed a unified network, termed as SiamCAT, to process HSIs that have varied numbers of bands. Besides, with a guided learning attention module, the spectral features can be highlighted by the spatial attention obtained from the RGB branch. Li et al. [28] designed a material-guided multiview fusion network MMF-Net, which combines the RGB, spectral, and material information to form a unified rich representation of an object. However, most existing methods simply adjust the number of input channels in the network to process HS data, resulting in limited spectral feature extraction. To address the above two issues, we construct a camouflaged object tracking dataset and propose a baseline model called SPDAN. SPDAN first utilizes a novel spectral embedding network to extract HS features. Then, a distractor-aware module is proposed to consider the position response from both the deep model and the Kalman filter, achieving robust tracking performance.
@@ II-B HOT Datasets
It can be seen that the above datasets neglect the camouflage characteristics of objects and only collect images in natural scenes, with most objects being vehicles and pedestrians, lacking the challenge of visual similarity. Although the WHU-Hi-H3 dataset considers the similarity between objects relative to the background, it does not highlight the visual confusion caused by the similarity between objects.
II-C Visual Object Tracker
@@
In recent years, the Transformer has been applied in many computer vision tasks such as detection [29], segmentation [30, 31], and multi-object tracking [32, 33, 34] due to its outstanding ability to model long-term dependencies. Many Transformer-based tracking methods have been proposed. Xie et al. [35] proposed a novel object-dependent feature network to embed the inter-image correlation. The non-object features can be suppressed in the self- and cross-attention schemes. Song et al. [36] designed a multi-scale cyclic shifting window attention to capture the window-level features. Mayer et al. [37] proposed a Transformer-based model prediction module, which captures global correlations without inductive bias. Cui et al. [38] presented a compact tracking framework MixFormer, which extracts and fuses features simultaneously through a mixed attention module without the need for additional fusion modules or box post-processing. Zhu et al. [20] developed a visual prompt tracking framework to achieve task-oriented multi-modality tracking. Zhou et al. [31] presented a detailed overview of the representative literature on both tracking methods and datasets. With the guidance of learned prompts, existing RGB-domain foundation models can effectively handle the downstream multi-modality tracking tasks. For HS trackers, there is currently limited research based on the Transformers.
II-D RGB/RGBT Tracking Datasets and Evaluation Metrics
The development of object tracking tasks has benefited from the promotion of a large number of publicly available datasets, particularly in the domains of RGB and RGB-T object tracking. RGB object tracking datasets are primarily used to evaluate tracking algorithms in visible conditions. Among them, the OTB series (OTB-2013[39], OTB-2015[40]) is one of the earliest benchmark datasets, encompassing various challenging factors such as object deformation, lighting variations, and occlusions. It employs success plots and precision plots as evaluation metrics. The VOT Challenge introduces new datasets annually, such as VOT2016, VOT2018, and VOT2019, which adopt strict evaluation criteria, including overlap accuracy and robustness. Additionally, they incorporate a re-initialization mechanism, making them critical benchmarks for evaluating tracking algorithms in competitive settings. In recent years, large-scale, long-term tracking datasets such as LaSOT[41] and TrackingNet[42] have further advanced the field. LaSOT consists of 1,400 long-duration video sequences, each averaging over 2,500 frames, making it a valuable resource for studying long-term object tracking. TrackingNet, on the other hand, is sourced from YouTube and contains over 30,000 video sequences, spanning a diverse range of real-world scenarios.
RGB-T object tracking datasets offer greater environmental adaptability by integrating RGB and thermal information. GTOT [43] is one of the earliest RGB-T tracking datasets, comprising 50 video sequences. The RGBT234 [44] dataset significantly expands this, containing 234 video sequences with a total frame count exceeding 234,000. It encompasses 12 challenging factors, including modality variations, camera motion, and object deformation, providing a more comprehensive benchmark for evaluating multi-modality object tracking algorithms. However, existing RGB-T datasets still face limitations in terms of diversity and scale. To address these shortcomings, the LasHeR [45] dataset has been introduced as the largest RGB-T object tracking benchmark. LasHeR consists of 1,224 video sequences with a total frame count exceeding 500,000, covering a wide range of scenarios such as indoor and outdoor environments, day and night conditions, and extended video durations, making it a crucial platform for studying long-term RGB-T tracking.
In recent years, many effective evaluation criteria have been proposed, including precision rate (PR), success rate (SR), accuracy, and Expected Average Overlap (EAO).
1) Precision rate is defined as the proportion of frames where the center location error (CLE) between the predicted bounding box and the ground truth remains below a specified threshold. The precision rate represents the percentage of such frames, while the precision plot illustrates how this rate varies as the threshold increases. Typically, a threshold of 20 pixels is used.
2) Success rate is determined by whether the overlap between the predicted bounding box and the ground truth exceeds a specified threshold. The success plot illustrates how the success rate varies as the threshold ranges from 0 to 1. To effectively compare different methods, the area under the curve (AUC) is used as a ranking metric.
3) Similar to the success rate, accuracy evaluates how well the predicted bounding box overlaps with the ground truth. It is determined by calculating the average overlap between the predicted and ground truth bounding boxes during successful tracking periods.
4) EAO systematically integrates per-frame accuracy and failure rates, providing a principled measure of a tracker’s expected no-reset overlap on short-term sequences. While other evaluation metrics exist, Precision Rate (PR) and Success Rate (SR) remain the most commonly used in HOT [14, 16], which is also what we use in our experiments.
@@ III-A Image Collection
We employ an IMEC XIMEA MQ022HG-IM-SM5X5-NIR camera and an EDMUND OPTICS 35mm C VIS-NIR lens to capture the high spatial and spectral resolution HS images across the wavelength range of 600-974 nm. This wavelength range includes a portion of the visible light spectrum. In this article, spectral information refers to the spectral information contained within the corresponding wavelength range of the dataset. The original image is a two-dimensional image with pixels of . After processing with the transformation algorithm, a 3D hyperspectral cube of can be obtained. To ensure the diversity of the scene, we collect data from different seasons and illuminations. The objects are real and artificial plastic fruits or vegetables in daily life to ensure the acquisition of camouflaged objects. which is shown in Fig. 2. We define camouflage in two aspects. First, camouflage occurs when the object and background share identical colors, similar to camouflage object detection in the RGB domain. Second, camouflage exists between different objects, primarily characterized by objects with the same appearance but different materials. To support research in this area, we acquire various objects in daily life, including plastic and metal bowls, cups, plastic fruits, and more. This allows us to create diverse object combinations and conduct experiments under background clutter and camouflaged object interference conditions. In addition, BihoT also provides false-color images obtained by selecting combinations of the 1st, 9th, and 15th bands, which serve as RGB modality in our experiments.
@@
III-B Data Annotation and Condition Challenges
The BihoT dataset has 49 annotated sequences, of which 27 are for training, and 22 are for testing. Each image has an object marked by a bounding box. All bounding boxes were manually created using the Labelme toolkit. In addition, we set 9 attributes corresponding to different challenging factors in the testing set, including Background Clutter (BC), Fast Motion (FM), In-Plane Rotation (IPR), Illumination Variation (IV), Low Resolution (LR), Occlusion (OCC), Out-of-Plane Rotation (OPR), Spectral Consistency (SC), and Spectral Variation (SV). Table II lists the number of bounding boxes and challenging scenarios for each sequence. To measure the spectral feature extraction ability of the tracker, we proposed two challenging factors, Spectral Consistency (SC) and Spectral Variation (SV), compared with the existing datasets [8, 10, 9]. Specifically, in the BihoT dataset settings, there are multiple real objects and one camouflaged object in each sequence. SC means that when the real object moves, its spectral information remains unchanged relative to other real objects in the scene. When the camouflaged object moves, the spectral information relative to other real objects in the scene is different. Therefore, the scene where the camouflaged object moves is denoted as SV.